# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ 6.2 Maximal Regularity of the Attractor-Completion of the Proof of Part II
$\|u(t)\|_{H^{2}}\leq R_{2},\quad\forall t\in\mathbb{R}.$
This concludes the proof of maximal regularity, hence of Part (II).
7 Appendix: Definitions and Some Relevant Results Pertaining to the Theory of Abstract Dynamical systems
The results of the abstract theory of attractors presented below can be found in [13, 14].
7.1 Basic Definitions
Let $(X,S_{t})$ be a dynamical system given by a $C_{0}$-semigroup $S_{t}$ on a complete metric space $X$.
Definition 7.1.
A bounded set $\mathfrak{B}\subset X$ is an absorbing set for $S_{t}$ if for any bounded set $\mathcal{B}\subset X$, there exists $t_{\mathcal{B}}\geq 0$ such that
$S_{t}\mathfrak{B}\subset\mathfrak{B},\quad\forall\,t\geq t_{\mathcal{B}},$
which defines $(X,S_{t})$ as a dissipative dynamical system.
Definition 7.2.
$(X,S_{t})$ is said to be asymptotically smooth in $X$ if for any bounded positive invariant set $\mathcal{B}\subset X$, there exists a compact set $K\subset\overline{\mathcal{B}}$, such that
$\mbox{dist}(S_{t}\mathcal{B},K)=0\quad\mbox{as}\quad t\to\infty.$
Definition 7.3.
A bounded closed set $\mathfrak{A}\subset X$ is said to be a global attractor of $(X,S_{t})$ iff the following properties hold: (i) $\mathfrak{A}\subset X$ is an invariant set, that is, $S_{t}\mathfrak{A}=\mathfrak{A}$ for all $t\geq 0$; (ii) $\mathfrak{A}\subset X$ is uniformly attracting, that is, for all bounded $D\subset X$
$\mbox{dist}(S_{t}D,\mathfrak{A})=\sup_{x\in S_{t}D}\inf_{y\in\mathfrak{A}}d(x,y)\to 0\quad\mbox{as}\quad t\to\infty.$
Definition 7.4.
Let $\mathcal{N}$ be the set of stationary points of $(X,S_{t})$:
$\mathcal{N}=\{v\in X:S_{t}v=v\;\;\mbox{for all}\;\;t\geq 0\}.$
We define the unstable manifold $\mathrm{M}^{u}(\mathcal{N})$ emanating from set $\mathcal{N}$ as a set of all $y\in X$ such that there exists a full trajectory $\Upsilon=\{u(t):t\in\mathbb{R}\}$ with the properties
$u(0)=y\quad\mbox{and}\quad\lim_{t\rightarrow-\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0.$
Definition 7.5.
$(X,S_{t})$ is said to be gradient if it possesses a strict Lyapunov function, i.e. there exists a continuous functional $\Phi(y)$ defined on $X$ such that
  • (i)
    the function $t\rightarrow\Phi(S_{t}y)$ is nonincreasing for any $y\in X$, and
  • (ii)
    the equation $\Phi(S_{t}y)=\Phi(y)$ for all $t>0$ and for some $y\in X$ implies that $S_{t}y=y$ for all $t>0$, i.e. $y$ is a stationary point of $(X,S_{t})$.
Definition 7.6.
The fractal dimension of a compact set $K\subset X$ is defined by
$\mbox{dim}K_{f}^{X}=\lim\sup_{\varepsilon\rightarrow 0}\frac{\ln(n(K,\varepsilon))}{\ln(1/\varepsilon)},$
where $n(X,\varepsilon)$ is the minimal number of closed balls of the radius $\varepsilon$ which cover the set $K$.
Definition 7.7.
A compact set $\mathfrak{A}_{\exp}\subset X$ is said to be inertial (or a fractal exponential attractor) of the dynamical system $(X,S_{t})$ iff $\mathfrak{A}_{\exp}$ is a positively invariant set of finite dimension and for every bounded set $D\subset X$ there exist positive constants $t_{D},C_{D}$ and $\gamma_{D}$ such that
$d_{X}\{S_{t}D|\mathfrak{A}_{\exp}\}\equiv\sup_{x\in D}\mbox{dist}_{X}\left(S_{t}x,\mathfrak{A}_{\exp}\right)\leq C_{D}e^{-\gamma_{D}(t-t_{D})},\quad t\geq t_{D}.$
Definition 7.8.
Let $X,Y$ be two reflexive Banach spaces with $X$ compactly embedded in $Y$ and set $H=X\times Y$. Consider the dynamical system $(H,S_{t})$ given by an evolution operator
$S_{t}z=(u(t),u_{t}(t)),\qquad z=(u_{0},u_{1})\in H,$
where the functions $u$ and $\theta$ possess the regularity
$u\in C(\mathbb{R}^{+};X)\cap C^{1}(\mathbb{R}^{+};Y),\quad$
Then one says that $(H,S_{t})$ is quasi-stable on a set $\mathcal{B}\subset H$ if there exist a compact seminorm $n_{X}$ on $X$ and nonnegative scalar functions $a(t)$ and $c(t)$ locally bounded in $0,\infty)$, and $b(t)\in L^{1}(\mathbb{R}^{+})$ with $\lim_{t\to\infty}b(t)=0,$ such that
$\|S(t)z^{1}-S(t)z^{2}\|_{H}^{2}$ $\leq$ $a(t)\|z^{1}-z^{2}\|_{H}^{2},$ (7.1)
$\|S(t)z^{1}-S(t)z^{2}\|_{H}^{2}$ $\leq$ $b(t)\|z^{1}-z^{2}\|_{H}^{2}+c(t)\sup_{0<s<t}\left[n_{X}(u^{1}(s)-u^{2}(s))\right]^{2},$ (7.2)
for any $z^{1},z^{2}\in\mathcal{B}$. The inequality (7.2) is often called quasi-stability inequality.
7.2 Abstract Results
We now present the statements of the abstract results used in the proof of Theorem 2.4. In order, the next theorem guarantees that a dissipative system has a compact global attractor if and only if it is asymptotically compact.
Theorem 7.1.
[14, Theorem 7.2.3] Any dissipative asymptotically smooth dynamical system $(X,S_{t})$ in a Banach space $X$ possesses a unique compact global attractor $\mathfrak{A}$. This attractor is connected set and can be described as set of all bounded full trajectories.
The following result gives the characterization of the attractor for gradient systems.
Theorem 7.2.
[14, Theorem 7.5.6] Let a dynamical system $(X,S_{t})$ possess a compact global attractor $\mathfrak{A}$. Assume that there exists a strict Lyapunov function on $\mathfrak{A}$. Then $\mathfrak{A}=\mathrm{M}^{u}(\mathcal{N})$. Moreover, the global attractor $\mathfrak{A}$ consists of full trajectories $\Upsilon=\{u(t):t\in\mathbb{R}\}$ with the properties
$\lim_{t\rightarrow-\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}\mbox{dist}_{X}(u(t),\mathcal{N})=0.$
The next three results show that quasi-stable systems enjoy nice properties that include both finite-dimensional, smoothness, and existence of exponential attractor.
Theorem 7.3.
[14, Theorem 7.9.6] Assume that the dynamical system $(H,S_{t})$ possess a compact global attractor $\mathfrak{A}$ and is quasi-stable on $\mathfrak{A}$. Then the attractor $\mathfrak{A}$ of has a finite fractal dimension $dim^{H}_{f}\mathfrak{A}$.
Theorem 7.4.
[14, Theorem 7.9.8] Assume that the dynamical system $(H,S_{t})$ possess a compact global attractor $\mathfrak{A}$ and is quasi-stable on $\mathfrak{A}$. Moreover, we assume that (7.2) holds with the function $c(t)$ possessing the property $c_{\infty}=\sup_{t\in\mathbb{R}^{+}}c(t)<\infty$. Then any full trajectory $\{(u(t);u_{t}(t);\theta(t)):t\in\mathbb{R}\}$ that belongs to the global attractor enjoys the following regularity properties,
$u_{t}\in L^{\infty}(\mathbb{R};X)\cap C(\mathbb{R};Y),\quad u_{tt}\in L^{\infty}(\mathbb{R};Y),\quad\theta\in L^{\infty}(\mathbb{R};Z).$
Moreover, there exists $R>0$ such that
$|u_{t}(t)|_{X}^{2}+|u_{tt}(t)|_{Y}^{2}+|\theta_{t}(t)|_{Z}^{2}\leq R^{2},\quad t\in\mathbb{R},$
where $R$ depends on the constant $c_{\infty}$, on the semigroup $\eta_{X}$ in Definition [14, Definition 7.9.2], also on the embedding properties of $X$ into $Y$.
Theorem 7.5.
[14, Theorem 7.9.9] Assume that the dynamical system $(H,S_{t})$ is dissipative and quasi-stable on some bounded absorbing set $\mathfrak{B}$. We also assume that there exists a space $\widetilde{H}\subset H$ such that $t\mapsto S_{t}z$ is Hölder continuous in $\widetilde{H}$ for every $z\in\mathfrak{B}$, that is, there exist $0<\gamma\leq 1$ and $C_{\mathfrak{B},T}>0$ such that
$|S_{t_{1}}z-S_{t_{2}}z|_{\widetilde{H}}\leq C_{\mathfrak{B},T}|t_{1}-t_{2}|^{\gamma},\quad t_{1},t_{2}\in[0,T],\;z\in\mathfrak{B}.$
Then, the dynamical system $(H,S_{t})$ possesses a (generalized) fractal exponential exponential attractor whose dimension is finite in the space $\widetilde{H}$.

### v4

@@ 6.2. Maximal Regularity of the Attractor-Completion of the Proof of Part II
7. Appendix-Ultimate Dissipativity
To conclude this work, we present the proof of the existence of an absorbing set for the system $(\mathcal{H},S_{t})$, which ensures that $(\mathcal{H},S_{t})$ is ultimately dissipative.
Proposition 7.1.
Proof.
Let $B$ be a bounded subset in $\mathcal{H}$ and let $U(t)=S_{t}U_{0}$ be a weak solution of problem (LABEL:P) with $U_{0}\in\mathcal{H}$. In order, let us consider the modified functional $\mathcal{J}(U(t)):=E(U(t))+L,$ where $L:=\frac{1}{\omega\lambda_{1}}\|h\|^{2}+C_{\nu}|\Omega|$ and $E(U(t))$ is defined in (2.9). Hence
[…]
Using the energy equality (2.8) with $s=t$ and $t=t+1$, we obtain
(7.2) $\int_{t}^{t+1}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(t))-E(U(t+1))=\mathcal{J}(U(t))-\mathcal{J}(U(t+1))=:\mathcal{D}(t).$
From Assumption (2.1), and and Hölder inequality with $\frac{2}{3}+\frac{1}{3}=1$, it follows from (7.2) that
(7.3) $\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq|\Omega|^{2/3}\left(\int_{t}^{t+1}\|u_{t}(s)\|^{6}_{6}ds\right)^{1/3}\leq\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}.$
(7.4) $\|u_{t}(t_{j})\|^{2}\leq 4\int_{t}^{t+1}\|u_{t}(s)\|^{2}ds\leq\frac{4|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3},\quad j=1,2.$
(7.5) $\begin{aligned} &\int_{t_{1}}^{t_{2}}E(s)ds+\frac{1}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds=\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds\\ &\;+\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx\;+\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\right]ds.\end{aligned}$
From Assumption (2.5), we get
(7.6) $\int_{t_{1}}^{t_{2}}\left[\int_{\Omega}F(u)dx-\int_{\Omega}f(u)udx\,\right]ds\leq\frac{\nu}{2}\int_{t_{1}}^{t_{2}}\int_{\Omega}|u|^{2}dxds\leq\frac{\nu}{2\lambda_{1}}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds.$
(7.7) $\begin{aligned} \int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds+\frac{\omega}{2}\int_{t_{1}}^{t_{2}}\|\nabla u(s)\|^{2}ds\leq&\;\frac{3}{2}\int_{t_{1}}^{t_{2}}\|u_{t}(s)\|^{2}ds-\underbrace{\int_{t_{1}}^{t_{2}}\int_{\Omega}g(u_{t})udxds}_{X_{1}}\\ &\;+\underbrace{\int_{\Omega}u(t_{1})u_{t}(t_{1})dx-\int_{\Omega}u(t_{2})u_{t}(t_{2})dx}_{X_{2}}+L.\end{aligned}$
$X_{1}$ $\leq$ $\kappa_{1}\left[\left(\int_{t}^{t+1}\|u_{t}\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|u\|^{2}ds\right)^{1/2}+\left(\int_{t_{1}}^{t_{2}}\|u_{t}\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|u\|^{6}_{6}ds\right)^{1/6}\right]$
$\leq$ $\frac{\kappa_{1}}{\lambda_{1}^{1/2}}\left[\left(\int_{t}^{t+1}\|u_{t}\|^{2}ds\right)^{1/2}\left(\int_{t}^{t+1}\|\nabla u\|^{2}ds\right)^{1/2}+\left(\int_{t_{1}}^{t_{2}}\|u_{t}\|^{6}_{6}ds\right)^{5/6}\left(\int_{t}^{t+1}\|\nabla u\|^{6}ds\right)^{1/6}\right]$
$\leq$ $\frac{2\kappa_{1}}{(\omega\lambda_{1})^{1/2}}\left[\left(\frac{|\Omega|^{2/3}}{\kappa_{0}^{1/3}}\mathcal{D}(t)^{1/3}\right)^{1/2}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}+\left(\frac{1}{\kappa_{0}}\mathcal{D}(t)\right)^{5/6}\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))^{1/2}\right]$
From embedding $H^{1}\hookrightarrow H^{0}$ and (7.4), we get
$X_{2}\leq\|u(t_{1})\|\|u_{t}(t_{1})\|+\|u(t_{2})\|\|u_{t}(t_{2})\|\leq\frac{16|\Omega|^{2/3}}{\varepsilon\omega\kappa_{0}^{1/3}\lambda_{1}}\mathcal{D}(t)^{1/3}+\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s)).$
Returning to (7.5), we obtain
(7.8) $\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds\leq C_{0}\mathcal{D}(t)^{1/3}+C_{1}\mathcal{D}(t)^{5/3}+2\varepsilon\sup_{t\leq s\leq t+1}\mathcal{J}(U(s))+L,$
(7.9) $\mathcal{J}(U(t+1))\leq\mathcal{J}(U(\tau))=\frac{1}{t_{2}-t_{1}}\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds\leq 2\int_{t_{1}}^{t_{2}}\mathcal{J}(U(s))ds.$
Hence, from (7.2), (7.9), and (7.8), we get
(7.10) $\mathcal{J}(U(t))^{3}\leq 4\left[\,4C_{0}+4C_{1}\mathcal{D}(t)^{4/3}+2\mathcal{D}(t)^{2/3}\,\right]^{3}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+[2^{8/3}L]^{3}.$
Using the definition of $\mathcal{D}(t)$ and the energy equality (2.8) we obtain that $|\mathcal{D}(t)|\leq 2|E(U_{0})|$. Hence,
Thus, (7.10) can be rewritten as
(7.11) $\mathcal{J}(U(t))^{3}\leq C_{0}\left[\,\mathcal{J}(U(t))-\mathcal{J}(U(t+1))\,\right]+\left[2^{8/3}L\right]^{3}.$
Applying Nakao’s lemma [29, Lemma 2.1] to (7.11), we conclude that
(7.12) $\mathcal{J}(U(t))\leq\left(2C_{0}^{-1}(t-1)^{+}+\mathcal{J}(U(0))^{-2}\right)^{-\frac{1}{2}}+2^{8/3}L.$
Using (7.1), it follows from (7.12) that
[…]